# L2M-Reg: Building-level Uncertainty-aware Registration of Outdoor LiDAR Point Clouds and Semantic 3D City Models

### v2

@@ 5.4 Limitations
As with most fine registration methods, L2M-Reg also requires an initial alignment. Developing a one-step solution that removes the dependency on coarse registration will be an important direction for future research.
Additionally, L2M-Reg assumes that the plinth regions are at least partially visible in the point clouds and not completely obstructed by vegetation or vehicles during data acquisition. Although a complete occlusion is very uncommon in practical scenarios, this assumption should be explicitly acknowledged.
@@ 6 Conclusion
This paper introduces L2M-Reg, a plane-based LiDAR-to-Model registration method tailored for individual buildings, with a particular focus on addressing the inherent uncertainty in LoD2 models. The key innovations of L2M-Reg are threefold. First, it considers uncertainty of LoD2 models, which serves as reference data, and introduces automated algorithms to identify representative regions and extract plane segments. Second, it introduces the concept of pseudo-planes in GHM and employs a 2D–3D decoupled parameter estimation strategy, which effectively mitigates the influence of low-reliability vertical model data on the estimation of horizontal parameters. Third, it maximizes the use of embedded semantic information in LoD2 models to enhance both the efficiency and reliability of plane correspondence establishment.
L2M-Reg was evaluated against leading solutions, including both ICP- and plane-based methods, across three real-world datasets. Experimental results show that L2M-Reg outperforms them in both accuracy and computational efficiency. Overall, L2M-Reg contributes toward bridging the gap in uncertainty-aware LiDAR-to-Model fine registration at the building level.
Accurate and efficient LiDAR-to-Model registration remains an open research challenge. Developing a one-step registration method that eliminates dependence on coarse registration represents a promising direction for future work. Additionally, extending the current L2M-Reg to support seamless registration across both indoor and outdoor environments presents another important avenue for further research.
@@
Ziyang Xu: Conceptualization, Methodology, Software, Validation, Investigation, Writing – review & editing, Writing – original draft, Visualization. Benedikt Schwab: Conceptualization, Resources, Data Curation, Writing - original draft, Writing – review & editing. Yihui Yang: Conceptualization, Validation, Visualization, Writing - original draft, Writing – review & editing. Thomas H. Kolbe: Supervision, Writing - Review & Editing. Christoph Holst: Supervision, Writing - Review & Editing, Funding acquisition, Project administration.
@@ Acknowledgments
This research was funded by TUM Georg Nemetschek Institute of Artificial Intelligence for the Built World, project ”NERF2BIM”, PI Christoph Holst. Many thanks to the City of Munich / GeodatenService for supporting this research.
@@
The datasets used in this paper and the implementation of code for L2M-Reg will be publicly available upon acceptance.

### v3

@@ 5.7 Limitations
Although this study employed exclusively real-world datasets for evaluation, given the complexity of real-world scenarios and other unresolved considerations, the current limitations are as follows:
  1. (1)
    As with most fine registration methods, L2M-Reg requires an initial alignment. Developing a one-step solution that removes the dependency on coarse registration will be an important direction for future research. L2M-Reg also assumes that the wall plinth areas are at least partially visible in the point clouds and not completely obstructed by vegetation or vehicles during data acquisition. Although a complete occlusion is very uncommon in practical scenarios, this assumption should be explicitly acknowledged.
  2. (2)
    When an entire facade is composed of transparent glass, most laser returns penetrate the surface and are recorded from interior structures, making it impossible to obtain reliable point clouds of the facade itself. In such cases, L2M-Reg is not applicable. This limitation is inherent to LiDAR-based point cloud acquisition.
  3. (3)
    As a plane-based method, L2M-Reg requires sufficient visibility of planar features in the input point clouds and becomes ineffective when such features are extremely sparse or nearly absent. If a facade is dominated by complex non-planar structures, such as geometrically irregular shapes, fully curved surfaces, arched structures, or wave-like forms, no dominant plane can be reliably identified. In these situations, the underlying assumption of the plane-based logic no longer holds.
  4. (4)
    A complete error propagation model was carefully considered. However, after detailed evaluation, it was determined that a rigorous error propagation analysis cannot be reliably provided at this stage for three main reasons. First, error sources in MLS point clouds are highly complex to model (Xu et al., 2025a, b). Second, the complete error modeling of TLS point clouds remains challenging (Holst and Kuhlmann, 2016). Third, the accuracy information provided by manufacturers may not be directly applicable. Given these unresolved issues, it is currently not feasible to construct a complete and reliable variance-covariance matrix for the point clouds used in this study. The investigation of the complete variance-covariance matrix is beyond the scope of this study. Relevant details can be found in these publications (Wujanz et al., 2017; Kerekes and Schwieger, 2020; AbdelGafar et al., 2025).
@@ 6 Conclusion
This paper introduces L2M-Reg, a plane-based LiDAR-to-Model registration method tailored for individual buildings, with a particular focus on addressing the inherent uncertainty in LoD2 models. The key innovations of L2M-Reg are threefold. First, it considers the uncertainties of LoD2 models that serve as reference data, and introduces automated algorithms to identify representative regions and extract plane segments. Second, it introduces the concept of pseudo-planes in GHM and employs a 2D-3D decoupled parameter estimation strategy, which effectively mitigates the influence of low-reliability vertical model data on the estimation of horizontal parameters. Third, it maximizes the use of embedded semantic information in LoD2 models to enhance both the efficiency and reliability of plane correspondence establishment.
L2M-Reg was evaluated against several leading solutions, including both ICP- and plane-based methods, across five real-world datasets. Overall, experimental results show that L2M-Reg outperforms them in both accuracy and computational efficiency. L2M-Reg contributes toward bridging the gap in uncertainty-aware LiDAR-to-Model fine registration at the building level. Moreover, this study enables accurate and cost-efficient model updating at the building level, extends the use of LoD2 models to accuracy-sensitive applications, and explicitly incorporates in-model uncertainty into downstream tasks. Together, this study advances both the methodological understanding and practical exploitation of uncertainty in geospatial data.
Accurate and efficient LiDAR-to-Model registration remains an open research challenge. Further application of L2M-Reg to a wider range of cities and larger-scale datasets, together with tailored optimizations, is planned. In addition, extending L2M-Reg to support seamless registration across both indoor and outdoor environments represents another important direction for future research.
@@
Ziyang Xu: Conceptualization, Data curation, Methodology, Software, Resources, Validation, Investigation, Writing – review & editing, Writing – original draft, Visualization. Benedikt Schwab: Conceptualization, Resources, Data curation, Writing - original draft, Writing – review & editing. Yihui Yang: Conceptualization, Validation, Visualization, Writing - original draft, Writing – review & editing. Thomas H. Kolbe: Supervision, Writing - Review & Editing. Christoph Holst: Supervision, Writing - Review & Editing, Funding acquisition, Project administration.
@@ Acknowledgments
This research was funded by TUM Georg Nemetschek Institute of Artificial Intelligence for the Built World, project ”NERF2BIM”, PI Christoph Holst. Many thanks to the City of Munich / GeodatenService for supporting this research, to 3D Mapping Solutions [^7] https://www.3d-mapping.de/ for providing the MLS point clouds of Ingolstadt as part of the SAVeNoW research project, and to Leonard Trill for data collection.
@@
The datasets used in this paper and the implementation of code for L2M-Reg can be found: https://github.com/Ziyang-Geodesy/L2M-Reg.